Tohban Report 2016-01-20
Tohban Reports | |
---|---|
Start Date: | 13 Jan 2016 |
End Date: | 20 Jan 2016 |
Tohban: | Jim McTiernan |
Tohban email: | jimm@ssl.berkeley.edu |
Next Tohban: | TBD |
List all reports |
Solar Activity
Not much,
How many GOES flares occurred?
Flares above B, C, M, X class were 27 4 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 3 1 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 3 1 0 0
There were RHESSI flares/GOES flares 9 / 31 over the time range 12-Jan-16 19-Jan-16
Memory Management
Decimation was active/vigorous for the full time period. Both front and rear detectors are off during SC night, except for 4 minutes on either side. The SSR was at 10% at the end of yesterday's passes.
On 15 January (day 015!) we skipped the playback pointer to free memory, this resulted in a vc3 data gap from 2016-014-23:30 to 2016-015-15:00 UTC.
Spacecraft Status
The cold plates, 1 and 2 were at 138.3 and 136.8 K, on 2016-01-18.
Data Gaps
No VC1 Gaps
VC3 Gaps GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30
GAP START TIME GAP END TIME GAP (SEC) 2016-01-14T15:35:00.000 -- 2016-01-14T15:40:00.000 300.00000 2016-01-14T23:20:00.000 -- 2016-01-15T15:05:00.000 57700.000
Detector issues
Detector 2 is unsegmented, Detector 9 is off, detectors 3, 4, 6 and 8 are not recording events. D6 was turned off beacuse it was filling the memory with noise.
D1 - HV was lowered to 2500 V
2016-013-21:20:47 /IHVDAC DETECTOR=1, VOLTAGE=129 ; 2500 V
D2 - No changes
D3 - No changes
D4 - On Jan 13, Detector 4 was lowered from 500 to 200 V. Events were recorded in sunlight for 3 minutes, and later for 8 minutes. The spectra seem ok, so we may want to try D4 for a full orbit:
2016-013-21:21:18 /IHVDAC DETECTOR=4, VOLTAGE=11 ; 200 V 2016-013-21:21:41 G4 events on 2016-013-21:24:39 G4 events off
2016-013-23:00:55 G4 events on 2016-013-23:08:06 G4 events off
For spectra see:
http://sprg.ssl.berkeley.edu/~jimm/temp/hsi_sepdet_spectrum_20160113_230000to230800.png
D5 - On 15 Jaunary we raised the Front Fast Threshold from 0x70 to 0x80, to increase live time; this reduced the Attenuator Average Deadtime to 17%, and popped out the thin attenuator, which had been in since approximately 22:00 on 14-Jan.
2016-015-01:46:23 start idib_chg_thrshld (5, FRONT, FAST, 0x80) 2016-015-01:46:27 /IDPUDUMPTABL TABLE=DIBTBL5 2016-015-01:46:27 /IDPUTABLE5 OFFSET=FRONTFASTDAC 2016-015-01:46:35 /IDPLOAD VALUE=0x80
D6 - On 13 January, we dropped the D6 HV to 500 V, to reduce noise and see if any spectrum could be recovered. No luck there, so on 15 january, we dropped the HV to 200 V. D6 desegmented and began to fill the memory. We raised the HV back to 500 V, and turned events off.
2016-013-23:02:09 /ihvdac detector=6, voltage=28 ; 500 V
2016-015-01:41:44 /ihvdac detector=6, voltage=11 2016-015-23:40:07 /ITMOFF VALUE=DIB6 ; off permanently 2016-015-23:40:25 /IHVDAC DETECTOR=6, VOLTAGE=28 ; 500 V
D7 - No changes
D8 - No changes
D9 - No changes
A spin up to 15 RPM was performed between 2016-013 21:23 and 22:59 UTC.
Spacecraft Management
Decimation | Active/Vigorous |
HLAT Decimation | Rear decimation weight 6, no front decimation |
Night time data (fronts) | +/- 4 minutes |
Night time data (rears) | +/- 4 minutes |
Require extra passes? | No |
Requirement for moving pointer? | Yes, see notes above |
Attenuator operation | Working as expected |
Detector problems? | See notes above. |